SOEST Home Search Site Map Links Contact
Dolphins at the bow of the Kilo Moana University of Hawaii Marine Center UH Marine Center Logo
Home Ships Marine Center Ship Techs Schedules Cruise Planning    

 

Overview

Layout

Handbook

Scientific Equipment

Support Equipment

Daily Report

Current Position

Schedule

Photo Album

 

Daily Report for R/V Kilo Moana July and August 2007


KM 07-15-07
Aug 29 2007

Greetings

DSL Survey continues, all going well

ETA "H" buoy 0715
ETA Pier 30 0745

Noon Pos. 21o12.9'N 157o55.2'W (5.6 nm. south of Honolulu)

55,431 gals F/O as 1200 hrs.
1043 gals consumed past 24 hrs.

wind 065o 13 kts.
swell SE 3'
brom. 1014.6 mlb.
temp. 82o F.
5/10 cloud cover



KM 07-15-06
Aug 28 2007

Greetings
DSL 120 survey continues off entrance to Honolulu and Pearl Harbors. we
are now on our most inshore survey line and will be crossing "PH" at 1317,
we then plan on moving offshore south of Barbers Pt Light. DSL is working
great and Science is happy with survey results so far.

Noon Pos. 21o16.0'N 157o56.6'W

56,474 gals. F/O on board as of 1200 hrs.
1,025 consumed past 24 hrs.

wind 5 kts all morning, just picked up to 105o 14 kts.
swell SE 3'
brom. 1014.6 mlb.
temp. 84o F.
4/10 cloud cover





KM 07-15-05
Aug 27 2007

Greetings one and all,

Deployed DSL 120 at 2041 (26th) and conducted survey off Pearl Harbor.
Recovered DSL 120 at 0956 this morning due to a electrical short on the
tow body connection. at this time it has been repaired and expect to
launch the DSL 120 around 1300 (+ or -)
all is going well, another beautiful day

Noon Pos. 21o14.7'N 158o00.4'W (8 nm. west of "H" Buoy)

57,499 gals. F/O on board as of 1200 hrs.
1,105 gals. consumed past 24 hrs.

wind 107o 15 kts.
swell SE 3'
brom. 1014.7 mlb.
83o F.
4/10 cloud cover

Gray


Sunday, August 26, 2007 12:37 PM
KM07-15-04

Greetings

we are continuing our survey south of Pearl Harbor with the 1002 system.
we expect to float test the DSL 120 this afternoon, and commence DSL 120
survey this evening or late afternoon.

Noon Pos. 21o08.4'N 158o00.1'W (10.5 nm. south of PH)

58,604 Gals. F/O on board as of 1200 hrs., consumed 1,192 Gals. past 24 hrs.

wind 124o 10 kts.
swell SE 2'
brom. 1014.7 mlb.
86o F.
3/10 cloud cover
all in all a very nice day

Gray



KM 07-15-02
24 August 2007

Greetings
Having a wonderful day at sea, watching the rain falling on Oahu this
morning, it seems to have cleared up this afternoon though.

We are conducting bottom survey between ten and twenty mile south of
Diamond Head, expect to be completed at 0500 25 Aug.

We still plan on picking up Steve T. Mark R.and Peter Dean of NAVTECH at
0800, by small boat at SNUG Saturday morning.

Noon Position 21o00.7'N 157o51.9'W
F/O on board at 1200hrs. 59,204 gallons (US)
Used 1,155 past 24 hrs.
wind 090 3-10 Kts. (nice)
Swell S'ly 1'
brom. 1015.3 mlb.
83o F.
5/10 cloud cover

All is going well and don't forget to stop, and enjoy your weekend.

Gray

 



Greetings
Sometimes its just hard to get out of town.
Ship whistle was repaired or should I say made operational by Ch.
Engineer, apparently there is a breaker I was not aware of, that is
labeled 3cm &10cm radar control panel, ships whistle control. I will
investigate how this breaker effects the ships radar as the radar was
fully functional when the breaker was open. Sorry for the delay
Last line 0923
"H" Buoy 1001
Noon Pos. 21o16.1'N 157o58.4'W
Fuel O/B 62,187

Wind 124o 15 Kts.
Swell 120o 2'
Brom. 1014.9 mlb.
86o F
6/10 cloud cover



KM 07-14-13
Monday Aug. 20 2007
KM 07-14-13
Greetings,

Medical evacuation of KM crew member in progress
Patient is doing much better and is resting at this time, providing
updates to MAS as requested.

1146 lost port propulsion, most likely due to overheating of cabinets in
switchboard romm, vessel was at 145 rpm with 4 eng. on line at time of
shut down. when restored may run at slower RPM. will keep you informed of
ETA Waianai small boat harbor.

Noon Position 22o 41.9'N 158o 25.1'W
Fuel report will follow
wind 080o 20 kts.
Swell E 6'
brom. 1018.0 mlb.
temp 82o F
4/10 cloud cover




KM 07-14-12

August 19 2007-1200hrs.

 

Noon position: 24o31.1’N 159o07.4’W

Fuel as of 1200hrs. 66,122 gals

Consumed 1,730 gals past 24 hrs.

 

WX: wind 075o  at 14Kts

Swell E 4ft.

Brom. 1021.0 mlb.

Sky 6/10 cloud cover

Temp 82oF

 

Recovered last array at 1315 this afternoon

 

Last night the bow thruster shut down, when we started it back up it would stay on for about 5 minutes and shut down again and again and again, we left it off all night and at 0747 I started it up, we did not use it but had it ready to go, it shut down once in a four hour period and when I used it to recover the array it shut down about 5 time within an hour.

The bow thruster will be required and operational for the next several trips, Roy Wilkens, Slow Speed Track lines and Jason Dynamic Positioning

Joe and John are in communication regarding Bow Thruster.

 

 

ETA “H” buoy 0700 Aug 21 2007

ETA Snug 0730-0745

 

Regards,

Gray

 


KM 07-14-11

August 18 2007-1200hrs.

 

Noon position: 23o06.4’N  159o37.1’W

Fuel as of 1200hrs. 67,852 gals

Consumed 1,267 gals past 24 hrs.

 

WX: wind 070o  at 15Kts

Swell E 4ft.

Brom. 1020.2 mlb.

Sky 6/10 cloud cover

Temp 83oF

 

Science work continues and slowly winding up. Recovered an array early this morning and will recover sediment trap, that was deploy over 7 days ago within the next hour, then steam northerly towards the second sediment trap which will be recovered tomorrow

All is well

 

John Nikola, Science wants to remove OTG van on starboard 01 level, I will come in Stbd. side too, with a few line, remove 20’ container, then shift to port side too and complete off load with ships crane.

 

ETA “H” buoy 0700 Aug 21 2007

ETA Snug 0730-0745

 

Regards,

Gray

 


KM 07-14-10

August 17 2007-1200hrs.

 

Noon position: 23o09.8’N  159o19.9’W

Fuel as of 1200 69,119 gals

Consumed 1,238 gals past 24 hrs.

 

WX: wind 080o  at 17Kts

Swell E 4ft.

Brom. 1020.0 mlb.

Sky 7/10 cloud cover

Temp 83oF

 

Science work continues, not much to report, other then all is going well with science crew and planned operations

 

Regards,

Gray

 


KM 07-14-09

August 16 2007-1200hrs.

 

Noon position: 23o12.8’N  159o08.6’W

Fuel as of 1200 70,357 gals

Consumed 1245 gals past 24 hrs.

 

WX: wind 085o  at 20Kts

Swell E 5ft.

Brom. 1018.3 mlb.

Sky 6/10 cloud cover

Temp 83oF

 

Science work continues, CTD and array deployments with an occasional Optics Cast, all is going well with Science.

 

10cm radar is down, will continue to correspond with NavTech for trouble shooting.

 

I repaired my computer myself, I was able to bring it up in safe mode, and under All Programs>Accessories> System Tools>System Restore>I was able to restore the computer as it was last Sunday Aug 12th.

 

I will try not to break any more equipment

Have a good weekend!!

 

Regards,

Gray

 


Greetings
KM 07-14-08
Wed. 15 Aug 2007

Noon Position 23o11.6'N  158o57.6'W (105 NM NNW Ka'ene Pt.)
Fuel oil as of 1200 Hrs. 71,602 Gals., used 2,603 Gals. past 24hrs.

Wx wind 085o 21kts.
Swell ESE 6'
Brom. 1018.0
4/10 cloud cover
Temp 84o F

Monitoring Tropical Storm Flossie!

Locating Sediment Trap with RDF and ARGOS positions and will be conducting
science ops in this area for the next few days, all going well, science
group is happy




KM 07-14-07

August 14 2007-1200hrs.

 

Noon position: 24o00.0’N 156o58.9’W

Fuel as of 1200 74,205 gals

Consumed 1523 gals

 

WX: wind 075o  at 20Kts

Swell E 8ft.

Brom. 1019.0 mlb.

Sky 6/10 cloud cover

Temp 81oF

 

conducted CTD work through the evening and morning All is going well with Science Ops

 

Continuing Monitoring and Plotting Tropical Storm FLOSSIE, Vessel moving towards the East, Conducting CTD operations every 10 nm along 24oN

As of 1230 we are steaming to 24oN and 156oW at 120 RPM for STA work. Then plan to make  way for Aloha Sta.

Regards,

Gray

 


KM 07-14-06

August 13 2007-1200hrs.

 

Noon position: 23o58.1’N 159o08.0’W

Fuel as of 1200 75,728 gals

Consumed 1348 gals

 

WX: wind 085o  at 20Kts

Swell ENE 6-7ft.

Brom. 1020.1 mlb.

Sky 2/10 cloud cover

Temp 83oF

 

Recovered Argon Gas Array this morning and conducted CTD work through the evening and morning All is going well with Science Ops

 

Monitoring and Plotting Tropical Storm FLOSSIE, Vessel moving towards the East, conducting CTD operations every 10 nm along 24oN as we cross area of  circulation , a area of interest for Sci.

Regards,

Gray

 


KM 07-14-05

August 12 2007-1200hrs.

 

Noon position: 23o51.5’N  159o14.4’W

Fuel as of 1200 77,076 gals

Consumed 1282 gals

 

WX: wind 100o  at 14 Kts

Swell ENE 5ft.

Brom. 1020.2 mlb.

Sky 8/10 cloud cover

Temp 79oF

 

On station conducting Science Ops, which include CTD, Optics, Trace Metal Cast’s and net tows, also deployed a Argon Gas Array this morning and recovered the Productivity Array last Night

All is going well with Science Ops

 

Monitoring and Plotting Tropical Storm FLOSSIE, looks like it may peter out in the next few days, but be assured I will not I would take my eyes off her

 

Regards

Gra


KM 07-14-02-04

 

August 11 2007-1200hrs.

 

Noon position: 23o46.7’N  159o13.6’W

Fuel as of 1200 78,358 gals

Consumed 1401 gals

 

WX: wind 095o  at 14 Kts

Swell 065o 4ft.

Brom. 1018.0 mlb.

Sky 7/10 cloud cover

Temp 83oF

 

On station conducting Science Ops, which include CTD, Optics, Trace Metal Cast’s and net tows, also deployed Sediment Traps and Productivity Array.

 

Monitoring and Plotting Tropical Storm FLOSSIE

 

Regards

Gray

 


KM 07-14-02-03

 

August 10 2007-1200hrs.

 

Noon position: 23o45.0N  159o14.6W (90NM North of Kaua’i)

 

Fuel as of 1200 79,759 gals

Consumed 1761 gals

 

WX: wind 075o  at 12 Kts

Swell 045o 3ft.

Brom. 1018.3 mlb.

Sky 6/10 cloud cover

Temp 84oF

 

0037 MVP Tow Fish lost due to parted cable. Evidence of shorted wire and arcing at point of separation.

 

Monitoring and Plotting Tropical Storm FLOSSIE, forecast at this time has storm passing 200 nm plus south of Honolulu

 

Regards

Gray

 


KM 07-14-02-02

 

August 9 2007-1200hrs.

 

Noon position: Ka’ena Pt. bearing 162o at 5.5nm.

 

Fuel as of 1200 81,520 gals

Consumed 537 gals

 

WX: wind 045o  at 22 Kts

Swell 045o 3ft.

Brom. 1016.0 mlb.

Sky 4/10 cloud cover

Temp 83oF

 

ETA Station 1645

 

Monitoring and Plotting Tropical Storm FLOSSIE

 

All is well with ship and crew

 

Regards

Gray

 


Greetings All,

 

KM-0713-04 Noon Report Saturday 04 August 2007

1200 Position 22o-47.8 N 158o-04.7 W

 

Fuel on board:    85,786 gal.

Fuel consumed:   1,113 gal.

 

The wind is ENE at 15 kts with a slight sea. The swell is NE at 5-6 ft. The baro is steady at 1018 MLB.

The sky is cloudy with 5/10’s cover, and the air temp is 81o F.

 

The gas array was recovered shortly after 0800 this morning. Station work continues at Aloha Station. I conferred with Susan Curliss, Chief Scientist concerning our propulsion problem. I explained to her that it was our intention to complete the station work at Aloha Station. It has been decided to drop the station off Kaena Point. Therefore, using a speed of advance of 7.0 kts we should make the “H” buoy at 1000 on Monday 06 August. I will keep you updated as our work here progresses.

 

Warmest Regards,

Phil  

 

 


Greetings All,

 

KM-0713-05 Noon Report Sunday 05 August 2007

1200 Position 22o-48.2 N 158o-12.2 W

 

Fuel on board:    84,556 gal.

Fuel consumed:   1,230 gal.

 

The wind is E’ly is 15 kts. with a slight sea. The swell is NE at 4-5 ft. The Baro is steady at 1018 MLB.

The sky is cloudy with 5/10’s cover, and the air temp is 82o F.

 

We recovered the sediment trap array which had drifted 32 nautical miles WNW. It was on board at 0905 this morning. I asked Joe about running back toward station using three generators and both shafts for a short period of time, to try to speed things up a bit. We managed to get 12 kts for just about one hour duration. The temps started rising again in the Stbd. Main Motor so we shut down again to single shaft operation. At 1315 we are back on station conducting science operations. I will confer with Susan Curliss later today to firm up a ETD from station. They were talking about doing a cast near the mooring which is 2 hours away from our present position. I will send a ETA message around 2200 after I see what kind of speed we can make realistically.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0713-03 Noon Report Friday 03 August 2007

1200 Position 22o-43.5 N 158o-00.5 W

 

Fuel on board:    86,899 gal.

Fuel consumed:   1,687 gal.

 

The wind is ENE at 18 kts with a moderate sea. The baro is steady at 1017 MLB.

The sky is partly cloudy with 4/10’s cover, and the air temp is 83o F.

 

Station work continues at Aloha Station. We are pretty much on schedule as far as science operations. We are maintaining station using the port main shaft and the bow thruster as needed. The Chief Engineer reports progress is being made in the change out of the internal blower in the Stbd. Main Motor. Hopefully we will be back on line with the stbd main shaft by early evening. I will send a message when all repairs have been completed…..

 

Warmest Regards,

Phil


Greetings All,

 

KM-0713-02  Noon Report Thursday 02 August 2007

1200 Position 21o-20.0 N 158o-15.7 W

 

Fuel on board:    88,586 gal.

Fuel consumed:      305 gal.

 

The wind is E’ly at 15 kts with a slight sea. The swell is ESE at 5-6 ft. The Baro is steady at 1016 MLB.

The sky is partly cloudy with 3/10’s cover, and the air temp is 83o F.

 

We arrived on Kahe Station at 1210 for a weight test cast prior to any CTD cast. The winch and crane operated flawlessly during the weight cast. Upon completion of Station Kahe we will proceed toward Aloha Station for HOT-194 station work.

 

Warmest Regards,

Phil



Greetings All,

 

KM-0712-02 Noon Report Monday 30 July 2007

1200 Position 22o-47.3 N 158o-05.1 W

 

Fuel on board:    89,124 gal

Fuel consumed:   1,924 gal.

 

The wind is ENE at 14 kts with a slight sea. The swell is NE at 5-6 ft. The Baro is steady at 1016 MLB.

The sky is partly cloudy with 2/10’s cover, the air temp is 86o F.

 

The recovery continues of the MOSEAN buoy. We are now hauling in the synthetic line portion of the mooring. The evolution started at 0755 with the rescue boat being launched to rig a recovery pendant to the buoy. After the pendant was attached the rescue boat rigged the hauling line to the float end of the array at 0825. The rescue was then recovered and secured for sea. The hard hat floats were recovered in a cluster at 0850. We should have the buoy on board by 1600. Our ETA for Pier 45 remains at 0930 tomorrow. We will wait for Wecoma to arrive then moor port side to. Being port side to will help facilitate off loading this group and load for the HOTS trip.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0711-03 Noon Report Wednesday 25 July 2007

1200 Position 18o-25.0 N 158o-17.9 W

 

Fuel on board:    96,623 gal.

Fuel consumed:   1,937 gal.

 

The wind is Ely at 22 kts with a moderate sea. The swell is ENE at 5-6 ft. The baro is steady at 1014 MLB.

The sky is cloudy with 6/10’s cover, and the air temp is 87o F.

 

Multi beam mapping goes on with developing and enhancing areas from previous surveyed areas along with new mapping on seamounts. We where able to shut down 1

Propulsion generator yesterday at 1730. We remained at 8.0 kts throughout the night until 0650 when a speed of 10.0 kts was required. We conducted our Quarterly Security Drill at 1230. It included a repel boarders drill scenario. The drill went on until being secured at 1310. All goes well with ship, crew and science.

 

Warmest Regards,

Phil   


Greetings All,

 

KM-0711-02 Noon Report Tuesday 24 July 2007

1200 Position 18o-46.8 N 158o-20.5 W

 

Fuel on board:    98,560 gal.

Fuel consumed:   2,526 gal.

 

The wind is ENE at 10 kts with a slight sea. The swell is E’ly at 5-6 ft. The Baro is dropping slowly at 1015 MLB.

The sky is cloudy with 6/10’s cover, and the air temp is 80o F.

 

The mapping goes on with good results. However three generators are required on line to maintain the speed of ten knots which is required at this point. There are other gaps from prior mapping cruises which may need to be further developed.

 

Warmest regards,

Phil

 


Greetings All,

 

KM-0711-01 Noon Report Monday 23 July 2007

1200 Position 20o-39.8 N 158o-09.5 W

 

Fuel on board:   100,289 gal.

Fuel consumed:       516 gal.

 

The wind is NE at 15 kts with a slight sea. The swell is ENE at 6-7 ft. The Baro is steady at 1017 MLB.

The sky is partly cloudy with 3/10’s cover, and the air temp is 81o F.

 

This report is late getting to you as I was involved in cleaning and dressing a crewmembers laceration to left wrist. All is well and there appears no further action will be required, save changing the dressing later today. The mapping commenced at 1315 at a speed of ten kts to start with. They have requested that needle gunning be kept at a minimum. Consequently the work which we had hoped to start, scaling on main deck aft port side will be have to wait for a more opportune time….

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-10 Noon Report Thursday 19 July 2007

1200 Position: 22o-13.3 N 159o-38.7 W

 

Fuel on board:    102,358 gal.

Fuel consumed:     1,436 gal.

 

The wind is ENE at 22 kts. with a moderate sea. The swell is NE at 5-6 ft. The Baro is steady at 1017 MLB.

The sky is partly cloudy with 4/10’s cover, and the air temp is 85o F.

 

We finished station work at 2000 last night and had way on for Kauai. We have been circumnavigating the Island of Kauai since 0730 this morning for the scientific contingent. At 1245 we were abeam of Haena Pt. 1.0 mile distant. I checked in with Pacific Missile Range at Barking Sands for submarine activity in the area. They stated there was no activity and we were free to enter the zone for transit. We will be at “H” buoy inbound at 0700 in the morning. We will see you alongside. Everything goes well with the ship and crew.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-09 Noon Report Wednesday 18 July 2007

1200 Position: 20o-43.4 N 159o-44.9 W

 

Fuel on board:    103,794 gal.

Fuel consumed:     1,410 gal.

 

The wind is ENE at 12 kts. with a slight sea. The swell is NE at 6-7 ft. The Baro is steady at 1018 MLB.

The sky is partly cloudy with 4/10’s cover, and the air temp is 83o F.

 

We arrived on eddy station Bigelow at 0100. We were approaching the sediment trap array at 1000. It was recovered in tact at 1040. At 1230 ISM training was conducted for the crew and scientific personnel. A practical demonstration on the proper use and discharge of parachute flares and floating orange smoke signals was given by Chief Mate and Third Mate.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-08 Noon Report Tuesday 17 July 2007

1200 Position: 20o-17.8 N 158o-58.8 W

 

Fuel on board:    105,204 gal.

Fuel consumed:     1,055 gal.

 

The wind is ESE at 10 kts with a slight sea. The swell is NE at 5-6 ft. The baro is rising slowly at 1017 MLB.

The sky is cloudy with 5/10’s cover, and the air temp is 81o F.

 

The sediment trap array at station Redfield was recovered at 1050 this morning. We are now proceeding to the North taking CTD’s every 15 miles toward eddy station Bigelow. Hurricane COSME has been downgraded to Tropical Storm status. It appears it will pass south of Oahu at a distance of 150 miles on the 21st.  All science operations are going well. All engineering and deck systems are up and running with no problems reported. The food and presentation have been excellent this entire voyage…as usual.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-07 Noon Report Monday 16 July 2007

1200 Position: 20o-18.2 N 158o-49.6 W

 

Fuel on board:    106,259 gal.

Fuel consumed:     2,462 gal.

 

The wind is ENE at 10 kts with a slight sea. The swell is Ely at 5-6 ft. The Baro is steady at 1016 MLB.

The sky is partly cloudy with 3/10’s cover, and the air temp is 82o F.

 

We arrived back on eddy station Redfield after steaming toward the east after departing Honolulu The satellite imagery gave indication of the eddy development site just as just South of the island of Lanai. We arrived around 2200, took a CTD cast then proceeded back to the WNW and eddy station Redfield. We were on station at 0800, and station work continues through out the day. We are keeping an eye on Tropical Storm COSME which was just elevated to Hurricane force status. It is supposed to pass just south of the Hawaiian Islands on the 21st of July.

 

Warmest Regards,

Phil

 


Greetings All,

 

KM-0710-06 Noon Report Sunday 15 July 2007

1200 Position 20o-53.6 N 158o-14.6 W

 

Fuel on board:   108,721 gal.

Fuel consumed:    1,416 gal.

 

The wind is NE at 18 kts. with a moderate sea. The swell is ENE at 6-7 ft. The Baro is steady at 1017 MLB.

The sky is partly cloudy with 3/10’s cover, and the air temp is 85o F.

 

We were hove to off Kalihi Channel Entrance at 0700. The rescue boat was launched and away at 0730 with M. Kinzie, D.Spurgin, Q. Perkins and K. Bjorkland scientist. The rescue boat was hoisted and on board at 0840 with boat crew, Bjorkland and Mike Rappe, Co-Chief Scientist. It was reported by boat crew that the throttle linkage to outboard motor is broken. Engineering dept. will look into the broken linkage. At 0850 we had way on toward eddy station Redfield. ETA for station Redfield is 1700 this afternoon.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-05 Noon Report Saturday 14 July 2007

1200 Position 20o-51.9 N 159o-14.7 W

 

Fuel on board:   110,137 gal.

Fuel consumed:    1,076 gal.

 

The wind is ENE at 11 kts. with a slight sea. The swell is ESE at 6-7 ft. The Baro is dropping slowly at 1017 MLB.

The sky is cloudy with 6/10’s cover, and the air temp is 81o F.

 

Station work continues on and near eddy station Bigelow. All science work is going well. All bridge and engine room monitors are up and running as before. We anticipate being off Kalihi Channel Entrance at 0700 tomorrow morning. The rescue boat will be launched around 0730 with Kinze, Spurgin and Perkins as boat crew carrying Karin Bjorkland as passenger. Karin needs to secure scientific supplies which are located at the Marine Center. Mike Rappe will join the vessel from Snug Harbor as well.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-04 Noon Report Friday 13 July 2007

1200 Position: 20o-45.7 N 158o-51.6 W

 

Fuel on board:   111,213 gal.

Fuel consumed:    1,439 gal.

 

The wind is NEly at 22 kts with a moderate sea. The swell is ESE at 6-7 ft. The Baro is steady at 1018 MLB.

The sky is partly cloudy with 4/10’s cover, and the air temp is 83o F.

 

We recovered the Productivity array last night at 1920  at station Redfield. The vessel is on station conducting a CTD cast enroute eddy station Bigelow. Last night the bridge secondary control computer was inadvertently reset and we are unable to reboot it to restore the current information. Steve Poulos suggested contacting Ross Ishida at RCF to see if he could be of assistance. The fact that unit is down has no bearing on vessel operation as the MCS computer in the engine room control station is working fine.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-03 Noon Report Thursday 12 July 2007

1200 Position: 20o-12.1 N 158o 33.2 W

 

Fuel on board:    112,652 gal.

Fuel consumed:     1,543 gal.

 

The wind is ENE at 10 kts with a slight sea. The swell is Ely at 6-7 ft. The Baro is steady at 1017 MLB.

The sky is partly cloudy with 3/10’s cover, and the air temp is 82o F.

 

Station work continues in and near the eddy Redfield. A productivity array was deployed at 0500 this morning for an 1800 recovery. All science operations are going well, with no reported problems or casualties. All engineering and deck machinery working with no problems reported. The galley staff is performing in an excellent fashion as always. It is certainly always nice to sail on board Kilo Moana.

 

Warmest Regards,

Phil


Greetings All,

 

KM-0710-02 Noon Report 1200 Wednesday 11 July 2007

 

Fuel on board:   114,195 gal.

Fuel consumed:    2,169 gal.

 

The wind is NEly at 12 kts. with a slight sea. The swell is ENE at 6-7 ft. The baro is steady at 1017 MLB.

The sky is partly cloudy with 4/10’s cover, and the air temp is 82o F.

 

After arrival on station last night we proceeded to map the area of the suspected eddy location. The first sediment trap array was deployed at 0620 at station # 1, Redfield. At the present time we are enroute station # 2, Bigelow. The crew was given their weekly ISM training as directed by Chief Mate. CTD winch and crane operation continue to go well since last weeks adjustments by Chief Engineer.

 

Warmest Regards,

Phil


Greetings All,

Km-0709-05 Noon Report 1200 Monday 09 July 2007

Fuel on board:   118,115 gal.

Fuel consumed:     1326 gal.

The wind is Ely at 15 kts with a slight sea. The swell is ESE at 6-7 ft. The Baro is steady at 1017 MLB.

The sky is partly cloudy with 3/10’s cover, and the air temp is 80o F.

The floating sediment trap array was recovered in tact at 0620. CTD casts and normal HOTS trip station work continues with no problems encountered. We will be conducting a pre-deployment meeting for all those who will be involved with the deployment this afternoon. This group will be deploying a permanently anchored sediment trap array for the first time. The Chief Mate and myself want to go over all aspects of this deployment with the participants. Our ETA remains 0700 at Hotel buoy and alongside portside to shortly after 0735. The OTG contingent desires to remove the HOTS tool van from main deck aft after this trip.

 

Warmest Regards,

Phil


Greetings All,

KM-0709-04 Noon Report 1200 Sunday 08 July 2007

1200 Position: 22o-43.6 N 158o-01-6 W

Fuel on board:   119,441 gal.

Fuel consumed:    1,121 gal.

The wind is ENE at 11 kts with a slight sea. The swell is E’ly at 4-6 ft. The Baro is steady at 1016 MLB.

The sky is cloudy with 6/10’s cover, and the air temp is 80o F.

Station work continues to progress smoothly. We have had no further problems with CTD winch since it was repaired by the Chief and 1st asst. engineer. The Chief Mate and myself conducted the monthly test of the SSAS, all in apparent good order.

Warmest Regards,

Phil


Greetings All,

KM-0709-02 Noon Report 1200 Friday 06 July 2007

1200 Position: 21o-20.6 N 158o-16.4 W

Fuel on board:   121,987 gal.

Fuel Consumed:      604 gal.

The wind is W’ly at 10 kts with a slight sea. The swell is SSE at 4-5 ft. The Baro is steady at 1019 MLB.

The sky is partly cloudy with 2/10’s cover, and the air temp. is 84o F. A really nice day out here at Kahe Station.

We arrived at Kahe Station shortly before noon. The hydro winch wire was lowered with a clump weight prior to any CTD casts. We anticipate staying at Kahe Station until around 1400 then proceed toward Aloha Station. All engineering systems are up and running with no apparent problems. Deck machinery; hydro winch and Appleton crane working as required, with no apparent problems as well.

Warmest Regards,

Phil



Archives - May and June 2007

Archives - March and April 2007

Archives - January and February 2007

 

 

 

 


School of Ocean and Earth Sciences and Technology, University of Hawaii

University of Hawaii Marine Center

1 Sand Island Access Road, Honolulu, HI 96819

Phone (808) 842-9813, Fax (808) 842-9833 Toll Free 1-888-800-0460

email snug@soest.hawaii.edu

University of Hawaii Link